OpenStudy (moongazer):

find the value of sin, cosine, tangent using these double angle identities. please check if my equation is correct. cos 2A = 1-2sin^2A then it will be

OpenStudy (moongazer):

\[sinA = \sqrt{(1-\cos2A)/2}\]

OpenStudy (moongazer):

and\[\cos2A = 2\cos ^{2}A-1\] will be\[cosA = \sqrt{(\cos2A+1)/2}\]

OpenStudy (asnaseer):

correct so far

OpenStudy (moongazer):

and \[tanA = sinA/cosA\] will be \[\sqrt{(1-\cos2A)/(\cos2A-1} = tanA\]

OpenStudy (asnaseer):

you got the last one wrong

OpenStudy (asnaseer):

denominator should have a plus sign instead of a minus sign

OpenStudy (asnaseer):

you seem to be pretty good at them already - just follow your instinct and use logical steps

OpenStudy (moongazer):

and do you have some tips to give so that will help me master these equations?

OpenStudy (asnaseer):

you don't need to remember all the equations. as long as you know the basic relationships, you should be able to derive all the rest.

OpenStudy (moongazer):

Actually, my teacher made the derivation I am just checking if I copied the correct one. :) But i'll derive it myself later so that I can practice. Thanks!

OpenStudy (asnaseer):

yw